OpenStudy (anonymous):

The length of a cube is 1/4 inch. What is the volume?

OpenStudy (solomonzelman):

All sides of the cube, are equal, so if the length of the cube is 1/4, then the height, and the width are also 1/4, each. That means that your volume, \(\color{#000000}{ \displaystyle V={\rm length}\times {\rm width}\times {\rm height} }\) \(\color{#000000}{ \displaystyle V=(1/4)\times (1/4)\times(1/4) }\)

OpenStudy (matthewrlee):

In what universe is the cube?